Raindrops are falling straight down at 12 m/s when suddenly the wind starts blowing horizontally at a brisk 4.0 m/s. From your point of view, the rain is now coming down at an angle. What is the angle? Express your answer using two significant figures.
Answer #1

tan theta = Y/x

tan theta = 12/4 = 3

theta = 72.16 deg

angle with horizontal = 90-72 = 18 deg